Grade 11Physical Chemistry

A vessel contains 12g of an ideal gas at 1°c temperature and 1 atm pressure.When the temperature is increased by 10°c at the same volume, the pressure increases by 10%. Calculate the volume and initial temperatute

sushmita
9 Years agoGrade 11

1 Answer

varun
9 Years ago
HELLO SUSHMITHA,
i will explain you this problem in two cases,
CASE 1
Given
P = 1 atm
W = 12g
T = (t + 273)K
V = V litre
 
CASE 2 
given 
        T = (t + 283)K
         P = 1+ 10/100
             = 1.1 atm
         W = 12 gm’
          V = V litre
Using the gas equation
 for case 1 = 1 X V  =  12/m X R(t + 273) -------------------(1)
 
for case 2 
                    1.1 X V = 12/m X R(t + 283) ------------------(2)
 
By equations (1) and (2)
 
1.1 / 1 = t + 283 / t + 273
therefore,
=> 1.1t + 300.3 = t  + 283
=> 0.1t = -17.3
=> t  = - 173C
 => t = 100K
and also from case 1 
since m = 120
1 X V = 12/120 X 100 X 0.082
=> V = 0.82L
THANK YOU